# Sample size and power calculation for propensity score analysis of observational studies

### v2

@@ 3.5 Analytical formula for sample size calculation
respectively, where the propensity score model parameters $(\mu_{e},\sigma_{e})$ and the outcome model parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ can be uniquely determined by the user specified inputs $(r,\phi)$ and $(r,\phi,E_{1},E_{0},S_{1},S_{0},R_{1},R_{0})$, respectively. The following theorem presents the analytical formula for the sample size calculation of the ATE (the proof is relegated to Appendix A.6).
@@
Let $\tau$ be the desired effect size of the ATE. The minimum sample size for a two-sided level-$\alpha$ test (4) for detecting $\tau$ with power $\beta$ is $N={V}\left(z_{1-\alpha/2}+z_{\beta}\right)^{2}/{\tau^{2}}$, where
$V=$ $\left(a_{1}^{2}+a_{0}^{2}\right)\sigma_{e}^{2}+\left(\sigma_{1}^{2}+\sigma_{0}^{2}\right)+\left[a_{1}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{1}^{2}\right]\exp\left(-\mu_{e}+{\sigma_{e}^{2}}/{2}\right)$
$+\left[a_{0}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{0}^{2}\right]\exp\left(\mu_{e}+{\sigma_{e}^{2}}/{2}\right).$ (19)
Remark 1. The above analytical derivation elucidates the critical role of covariate overlap in sample size calculations of observational studies. Specifically, the overlap coefficient $\phi$ explicitly determines the propensity score distribution, which in turn determines the potential outcome distributions because of the projection of $Y(z)$ on $W_{e}$. Intuitively, as the overlap decreases, the observational study deviates further from a target trial, rendering the marginal distribution of the observed outcomes deviate further from that of the potential outcomes. Therefore, to recover causal estimands, which are functions of potential outcomes, more sample units are required in observational studies with limited overlap.
@@
Remark 2. For binary outcomes, because the mean of a Bernoulli distribution determines the variance, compared to the continuous outcomes in Section 3.3, we have two fewer constraints to determine the outcome distribution. As a simple solution, we recommend to use linear models for binary outcomes so that the proposed methodology directly applies. Simulations (Section 5.2) and real applications (Section 6) later suggest that such a method generally works well in practice.
Remark 3. IPW is prone to extreme weights, which lead to inflated variance of the Hájek estimator and thus much reduced power. The common remedy is trimming, i.e. to exclusive units with extreme weights (Crump et al.,, 2009). However, trimming reduces the actual sample size, and more importantly, changes the target estimand to the population with substantial overlap. When extreme weights are expected, we recommend the users to consider alternative estimands targeting at the overlapped population, such as the ATO, and apply the related sample size calculation formulas developed in Section 4.
4 Weighted average treatment effect estimands
$\tau^{h}=\frac{\operatorname{\mathbb{E}}[\tau(x)h(x)]}{\operatorname{\mathbb{E}}[h(x)]}.$ (20)
The Hájek estimator for the WATE is:
$\widehat{\tau}_{w}=\frac{\sum_{i}w_{1}(X_{i})Z_{i}Y_{i}}{\sum_{i}w_{1}(X_{i})Z_{i}}-\frac{\sum_{i}w_{0}(X_{i})(1-Z_{i})Y_{i}}{\sum_{i}w_{0}(X_{i})(1-Z_{i})}.$ (21)
$V_{w}\equiv\operatorname{\mathbb{V}}(\widehat{\tau}_{w})=\frac{1}{\mathbb{E}[h(X_{i})^{2}]}\mathbb{E}\left[\left\{\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}+\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}\right\}h(X_{i})^{2}\right],$ (22)
$\int_{-\infty}^{\infty}h(u)^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u,$ (23)
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$ $=a_{1}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (24)
$\mathbb{E}\left[\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}h(X_{i})^{2}\right]$ $=a_{0}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (25)
5 Simulation
5.1 Simulation design
@@
Following the simulation design in Li et al., (2019), we simulate $N=1000$ units with 10 covariates of different types. Covariates $X_{1}$ to $X_{4}$ are binary with $\mathsf{Ber}(p)$ with $p=0.2,0.4,0.6,0.8$ respectively; $X_{5}\sim\mathsf{Unif}(0,1)$; $X_{6}$ to $X_{8}$ follows a Poisson distribution with mean parameters $1,2,$ and $3$ respectively; $X_{9}\sim\mathsf{Gamma}(2,3)$ and $X_{10}\sim\mathsf{Beta}(2,3)$. We generate the treatment indicator $Z\sim\mathsf{Ber}(\operatorname{\mathrm{e}xpit}(\beta_{0}+\kappa X\beta))$, where $\beta=(1,1,-1,0,-2,1,0.5,0,0,0)^{\prime}$ and $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$ is the parameter that controls the overlap between two treatment groups. For these $\alpha$ values, we choose $\beta_{0}=\{0,-0.248,-0.489,-0.722,-0.860,-0.951\}$ accordingly so that the treatment proportion $r=\mathbb{E}[Z]=0.5$. We then simulate the potential outcome $Y(z)\sim\mathsf{N}(X\gamma+\tau z,4^{2})$, where $\gamma=(1,1,-1,-1,0,-1,-1,0,1,1)^{\prime}$, and let the observed outcome $Y_{i}=Z_{i}Y(1)+(1-Z_{i})Y(0)$. The outcome model implies a homogeneous treatment effect $\tau=1$ and thus the true ATE, ATT, ATO are all $1$.
The overlap between two treatment groups is shown in Figure 2: $\kappa=0$ corresponds to randomized trials; as $\kappa$ increases, the overlap decreases. For each of the six simulation settings, we simulate $B=1000$ datasets and calculate the ATE of each dataset using the Hájek estimator (2) with a $95\%$ confidence interval, with both the true and estimated propensity scores. We then approximate the power by the proportion of the datasets where the confidence interval does not cover zero. For $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$, which corresponds to the overlap coefficient $\phi\in\{1.00,0.98,0.93,0.87,0.84,0.81\}$, respectively, the power with the true propensity score is $0.944,0.931,0.896,0.788,0.683,0.612$, respectively, and the power with the estimated propensity score is $0.978,0.972,0.943,0.852,0.737,0.659$, respectively. As expected, with the same sample size, decreased degree of overlap leads to much lower power.
@@
Now we fix the calculated power as the desired level of power, and inversely solve for the sample size, which should be close to $1000$. For a given simulated dataset $(X,Y,Z)$, we calculate the eight summary statistics $(r,\phi,E_{1},E_{0},S_{1},S_{0},R_{1},R_{0})$ as discussed in Section 3. The proportion of treatment, $r$, is calculated by the sample mean $\overline{Z}$, and the observed mean and variance, $E_{1},E_{0},S_{1},S_{0}$ are calculated with values of $Y$ on subsets $Z=1,0$, respectively. With the logistic model for the propensity score, $e(X)=\operatorname{\mathrm{e}xpit}(\beta_{0}+X\beta)$, the latent variable $W_{e}$ is estimated by $\widehat{\beta}_{0}+X\widehat{\beta}$, where $(\widehat{\beta}_{0},\widehat{\beta})$ is the estimated intercept and coefficients of the logistic regression of $Z$ on $X$. Although none of the covariates $X$ is normal, the latent linear combination of $X$, $W_{e}$, appears to be close to normal (see the rightmost column of Figure 4 in Appendix B.1). The calculation of $\phi$ is more complicated, and we calculate it using the empirical CDF of the propensity scores (details are relegated to Appendix B.3). We present $(r,\phi)$ in Table 1 below, and the remaining quantities $(E_{1},E_{0},S_{1},S_{0},R_{1},R_{0})$ are shown in Table 3 in Appendix C.2.
Continuous outcome Binary outcome
True PS Estimated PS True PS Estimated PS
$\phi$ Proposed $z$-test Proposed $z$-test Proposed $z$-test Proposed $z$-test
1.00 1005 1009 1266 1275 977 977 1119 1119
0.98 992 951 1244 1202 1016 973 1162 1114
0.93 1003 827 1210 998 1036 867 1168 980
0.88 979 609 1162 719 1079 680 1153 737
0.84 980 467 1140 538 1115 543 1148 575
0.81 1065 398 1179 445 1174 460 1307 518
@@
Table 1: Sample size calculation by the proposed method and the two-sample $z$-test to achieve nominal power under various degrees of overlap. The nominal power in each row is the true power of Hájek estimator for a sample of $N=1000$ units, with true and estimated propensity scores, respectively.
The left side of Table 1 shows the calculated sample size to achieve the corresponding power under various degrees of overlap. The sample size to achieve nominal power of the ATE calculated with the true propensity score should theoretically be 1000, which is very close to our calculated results when overlap is good. As the overlap decreases, the discrepancy between the calculated and the theoretical sample size increases, albeit still within a reasonable range. As noted before, the sample size calculated based on the estimated propensity scores is conservative, i.e. larger than the nominal size. The simulation results confirm this: the calculated sample size is slightly larger than the true sample size as long as the two treatment groups have a modest overlap. We also include the sample size calculated by two-sample $z$-test for comparison. When the overlap is perfect, the $z$-test correctly calculates the sample size. However, when the overlap is poor, meaning that the intended observational study is far from a target randomized trial, the $z$-test, which treats observational studies as randomized trials and thus ignore the degree of overlap, tends to vastly underestimate the sample size. Another observation is that the ratio between the $z$-test based sample size and the nominal counterpart varies by the degree of overlap. Therefore, the existing approach of inflating the $z$-test based sample size with a fixed variance inflation factor (e.g. Austin,, 2021) would lead to inaccurate sample sizes.
5.2 Binary outcome
We repeat the simulation design as in Section 5.1, and dichotomize the potential outcome $Y_{i}(z)$ into a binary variable $Y^{b}_{i}(z)=\mathbb{I}(Y_{i}(z)>-2)$. By SUTVA, the observed outcome $Y^{b}_{i}=Z_{i}Y^{b}_{i}(1)+(1-Z_{i})Y^{b}_{i}(0)$. The true ATE on $Y^{b}$ is $\tau^{b}=\mathbb{E}[Y^{b}_{i}(1)-Y^{b}_{i}(0)]=0.089$.
Assume we estimate $\tau^{b}$ using the Hájek estimator with $Y^{b}_{i}$. For $\kappa=\{0,0.25,0.5,0.75,0.9,1\}$, the power with the estimated propensity scores is $0.849,0.846,0.797,0.677,0.570,0.527$, respectively, and the power with the true propensity scores is $0.798,0.795,0.747,0.642,0.546,0.480$, respectively. Dichotomizing the outcome does not affect the propensity scores, and hence $(r,\phi)$ remain the same as in that in Simulation 5.1. We inversely solve for the sample size, and show the results on the right side of Table 1.
The results show that the proposed method, regardless of whether the true or estimated propensity scores are used, yields sample sizes close to the nominal value $1,000$. In contrast, the two-sample $z$-test continues to lead to vastly underestimated sample size as the degree of overlap decreases. Unlike the continuous outcome case, when the outcome is binary we tend to overestimate the sample size as the overlap coefficient decreases.
5.3 ATT and ATO estimands
With the same simulation design and data generating process as in Section 5.1, we focus on the ATT and ATO estimands, the true value of which are both 1. For $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$, corresponding to $\phi\in\{1.00,0.98,0.93,0.87,0.84,0.81\}$, we calculate the power of the Hájek estimator (21) for the ATT and ATO, with estimated and true propensity scores, respectively. Then we inverse the process to find the sample size needed under each scenario to obtain the corresponding nominal power; the results are shown in Table 2. In all scenarios, the estimated sample size is close to the true value 1000. As predicted from the general theory on the WATE (Li et al.,, 2018), the ATT estimator has large variance than the ATO estimator. Consequently, given the same setting, the ATT requires larger sample size to achieve the same level of power the ATO, and the discrepancy increases as the degree of overlap decreases.
ATT ATO
True PS Estimated PS True PS Estimated PS
$\phi$ Power Size Power Size Power Size Power Size
1.00 0.956 1083 0.969 1178 0.936 971 0.969 1184
0.98 0.910 919 0.966 1197 0.934 988 0.966 1194
0.93 0.837 842 0.897 1026 0.920 1006 0.972 1330
0.88 0.717 854 0.711 854 0.889 959 0.923 1082
0.84 0.632 965 0.641 975 0.886 974 0.916 1085
0.81 0.593 1124 0.546 1019 0.859 912 0.904 1052
Table 2: Sample size (size) calculated by the proposed method to achieve the nominal power (power) under $r=0.5$ and various degrees of overlap in simulations. The nominal power is the true power of Hájek estimator for a sample of $N=1000$ units, with the true and estimated propensity scores, respectively.
As noted in Section 4, the sample size for the ATT and ATO calculated using the estimated propensity scores is not guaranteed to be conservative. However, this simulation shows that (i) the calculated sample sizes under the true and estimated propensity scores are similar, and (ii) the sample size calculated for the ATO appears to be conservative if one expects to use the estimated propensity score in the study.
6 Real application: Right heart catheterization study

### v3

@@ 3.5 Analytical formula for sample size calculation
respectively, where the propensity score model parameters $(\mu_{e},\sigma_{e})$ and the outcome model parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ can be uniquely determined by the user specified inputs $(r,\phi)$ and $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$, respectively. To further reduce the number of inputs, we impose $S_{1}^{2}=S_{0}^{2}=S^{2}$ and $\rho_{1}=\rho_{0}=\rho$. Empirical evidence from multiple real data, e.g. the ones investigated in the supplementary materials, supported this restriction. Furthermore, following the convention in previous power analysis, we define the standardized effect size as the treatment effect scaled by the standard error: $\widetilde{\tau}=\tau/S$.
The following theorem presents the analytical formula for the sample size calculation of the $\tau^{\mbox{\tiny{ATE}}}$ (the proof is relegated to Supplementary Material A.6).
@@
Let $\widetilde{\tau}$ be the desired standardized effect size of the ATE. The minimum sample size for a two-sided level-$\alpha$ test (4) for detecting $\widetilde{\tau}$ with power $\beta$ is $N=\widetilde{V}\left(z_{1-\alpha/2}+z_{\beta}\right)^{2}/{\widetilde{\tau}^{2}}$, with
$\widetilde{V}=$ $\rho^{2}\left(\frac{\sigma_{e}^{2}}{\sigma_{e\mid 1}^{2}}+\frac{\sigma_{e}^{2}}{\sigma_{e\mid 0}^{2}}\right)+2(1-\rho^{2})+\left\{\rho^{2}\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 1}^{2}}+(1-\rho^{2})\right\}\exp(-\mu_{e}+\sigma_{e}^{2}/2)$
+$+\left\{\rho^{2}\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 0}^{2}}+(1-\rho^{2})\right\}\exp(\mu_{e}+\sigma_{e}^{2}/2),$ (17)
where $\sigma_{e\mid z}^{2}=\mathbb{V}(W_{e}\mid Z=z)$.
Theorem 2 shows that the formula of $\widetilde{V}$ does not depend on $\mu_{z}$. Moreover, equation (12) shows that $E_{z}$ is involved in solving $\mu_{z}$ but not $a_{z}$ and $\sigma_{z}^{2}$. Therefore, calculating $\widetilde{V}$ does not require $E_{z}$, further reducing the number of inputs. Indeed, the formula of $\widetilde{V}$ only involves $\sigma_{e}^{2}$, $\mu_{e}$ and $\sigma_{e\mid z}^{2}$, all of which can be derived from the distribution of $e(X)$ that is determined by $(r,\phi)$, and an additional input $\rho$. This simplification reduces the number of inputs for the sample size calculation to two, namely $(\phi,\rho)$, in addition to the standard $\widetilde{\tau},\alpha,\beta,r$. In the special case of a randomized trial, $\rho=0$, and the minimal sample size reduces to formula (6).
Remark 1. The above analytical derivation elucidates the connection and distinction between sample size calculations in randomized trials and observational studies. Specifically, four parameters — $\alpha$, $\beta$, the standardized effect size $\widetilde{\tau}$, and the treatment proportion $r$—are required in both cases. For observational studies, two additional inputs are required: the overlap coefficient $\phi$ and the correlation parameter $\rho^{2}$, both of which have natural interpretation. This result is intuitive because the key difference between observational studies and randomized trials is the presence of confounders in the former, which are pre-treatment covariates that affect the treatment assignment and the outcome simultaneously. The parameters $\phi$ and $\rho^{2}$ quantify the strength of confounder-treatment and confounder-outcome correlation, respectively. Similar strategy is used in the formulation of the E-value (Ding and VanderWeele,, 2016). Intuitively, as the overlap decreases (i.e., $\phi$ deceases) and the correlation between covariates and the outcome increases (i.e., $\rho^{2}$ increases), the observational study deviates further from a randomized trial, rendering the marginal distribution of the observed outcomes deviate further from that of the potential outcomes and thus require more sample units to recover the causal estimands. Indeed, the calculated sample size monotonically increases with $\rho^{2}$ from (17) (the proof is relegated to Supplementary Material A.7.). Moreover, numerical experiments show that the calculated sample size monotonically decreases with $\phi$.
Remark 2. The sample size formula of Shook-Sa and Hudgens, (2022) – derived based on the variance inflation factor – is a special case of formula (17). Specifically, when assuming the propensity score distribution $e(X)\sim\operatorname{\mathrm{e}xpit}(\mathsf{N}(\mu_{e},\sigma_{e}^{2}))$, their formula becomes $\widetilde{V}_{\mathrm{SH}}=2+\exp(-\mu_{e}+\sigma_{e}^{2}/2)+\exp(\mu_{e}+\sigma_{e}^{2}/2)$, which is formula (17) with $\rho=0$. We show that variance $\widetilde{V}_{\mathrm{SH}}$ is smaller than the variance $\widetilde{V}$ in (17) by a term proportional to $\rho^{2}$ (details in Supplementary Material A.7). Although $\rho^{2}$ is small in many of our simulation studies, ignoring it risks underestimating the sample size in general settings. Therefore, we recommend to report power calculations with a range of $\rho^{2}$ values between 0 and $R^{2}_{Y|X}$. Shook-Sa and Hudgens, (2022) discussed a remainder term that implicitly adjusts for the confounder-treatment association, but did not point out the connection, and the remainder term is only estimable with prior data on both weights and outcomes. In contrast, our method does not require this information and explicitly formulates the role of the confounder-treatment association.
@@
Remark 3. For binary outcomes, because the mean of a Bernoulli distribution determines the variance, compared to the continuous outcomes in Section 3.3, we have two fewer constraints to determine the outcome distribution. As a simple solution, we recommend to use linear models for binary outcomes so that the proposed method directly applies. Numerical examples in Section 4.2 and Supplementary Material D.3 suggest that this approach generally works well in practice.
4 Numerical Examples
4.1 Synthetic Data
@@
Following the simulation design in Li et al., (2019), we simulate $N=1,000,000$ units with 10 covariates of different types as the superpopulation. Covariates $X_{1}$ to $X_{4}$ are binary with $\mathsf{Ber}(p)$ with $p=0.2,0.4,0.6,0.8$ respectively; $X_{5}\sim\mathsf{Unif}(0,1)$; $X_{6}$ to $X_{8}$ follows a Poisson distribution with mean parameters $1,2,$ and $3$ respectively; $X_{9}\sim\mathsf{Gamma}(2,3)$ and $X_{10}\sim\mathsf{Beta}(2,3)$. We generate the treatment indicator $Z\sim\mathsf{Ber}(\operatorname{\mathrm{e}xpit}(\beta_{0}+\kappa X\beta))$, where $\beta=(1,1,-1,0,-2,1,0.5,0,0,0)^{\prime}$ and $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$ is the parameter that controls the overlap between two treatment groups. For these $\alpha$ values, we choose $\beta_{0}=\{0,-0.248,-0.489,-0.722,-0.860,-0.951\}$ accordingly so that the treatment proportion $r=\mathbb{E}(Z)=0.5$. We then simulate the potential outcome $Y(z)\sim\mathsf{N}(X\gamma+\tau z,4^{2})$, where $\gamma=(1,1,-1,-1,0,-1,-1,0,1,1)^{\prime}$, and let the observed outcome $Y_{i}=Z_{i}Y(1)+(1-Z_{i})Y(0)$. The outcome model implies a homogeneous treatment effect $\tau(x)=1$ and thus all average effects including $\tau^{\mbox{\tiny{ATE}}}$ equal 1.
The overlap between two treatment groups is shown in Figure 2: $\kappa=0$ corresponds to randomized trials; as $\kappa$ increases, the overlap decreases. The values of $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$ correspond to the overlap coefficient $\phi\in\{1.00,0.98,0.93,0.87,0.84,0.81\}$, respectively.
@@
Now we fix the power at 0.80, and solve for the sample size under each value of $\kappa$. We calculate the eight summary statistics $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$ as discussed in Section 3 from the $N=1,000,000$ units in the superpopulation. The proportion of treatment, $r$, is calculated by the sample mean $\overline{Z}$, and the observed mean and variance, $E_{1},E_{0},S_{1}^{2},S_{0}^{2}$ are calculated with values of $Y$ on subsets $Z=1,0$, respectively. With the logistic model for the propensity score, $e(X)=\operatorname{\mathrm{e}xpit}(\beta_{0}+X\beta)$, the latent variable $W_{e}$ is estimated by $\widehat{\beta}_{0}+X\widehat{\beta}$, where $(\widehat{\beta}_{0},\widehat{\beta})$ is the estimated intercept and coefficients of the logistic regression of $Z$ on $X$. Although none of the covariates $X$ is normal, the latent linear combination of $X$, $W_{e}$, appears to be close to normal (see the rightmost column of Figure 4 in Supplementary Material B.1), and the correlation parameters $\rho_{z}$ can be estimated by the sample correlation between estimated $W_{e}$ and $Y(z)$. We also calculate the pooled variance $S^{2}$ and pooled correlation $\rho$ empirically. To calculate $\phi$, we use the empirical cumulative probability function of the propensity scores (details are relegated to Supplementary Material B.3). We present $(r,\phi)$ in Table 1 below, and the remaining quantities $(E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$ and the pooled quantities $(S^{2},\rho^{2})$ are shown in Table 2 in Supplementary Material C.2. We also calculate and show in Table 2 the R-squared statistic $R^{2}$ of linear model $Y\sim X$, with $Y$ centered within the respective treatment arm. The results show that $S_{1}^{2}\approx S_{0}^{2}\approx S^{2}$ and $\rho_{1}^{2}\approx\rho_{0}^{2}\approx\rho^{2}$ in all settings. The $R^{2}$ turns out to be much larger than $\rho^{2}$, which is not surprising as $R^{2}$ is the theoretical upper bound of $\rho^{2}$, and provides the most conservative estimate when there is no substantive knowledge on the range of $\rho^{2}$.
Proposed $z$-test Shook-sa & Hudgens
$\phi$ Size True PS Est PS Size True PS Est PS Size True PS Est PS
1.00 645 0.83 0.88 629 0.80 0.87 629 0.80 0.87
0.98 656 0.79 0.87 628 0.78 0.85 653 0.80 0.88
0.93 754 0.79 0.87 626 0.73 0.80 742 0.78 0.86
0.88 993 0.78 0.85 624 0.60 0.65 950 0.77 0.84
0.84 1290 0.80 0.85 623 0.48 0.53 1216 0.77 0.84
0.81 1617 0.81 0.87 622 0.40 0.45 1481 0.79 0.81
@@
Table 1: Sample size calculation by the proposed method, the two-sample $z$-test and the method in (Shook-Sa and Hudgens,, 2022) to achieve nominal power of 0.80 under various degrees of overlap. The power is obtained from random sampling of the calculated size from the $N=1,000,000$ units. The Monte Carlo error for a power of 0.8 is 0.004.
We use inputs $(r,\phi,\rho^{2})$ and standardized effect size $\widetilde{\tau}=1/S$ to inversely calculate the sample size $N_{\mathrm{pos}}$. Table 1 shows the calculated sample size to achieve the nominal power 0.80 under various degrees of overlap. Then, we sample $N_{\mathrm{pos}}$ units randomly from the $N=1,000,000$ units in the superpopulation. We repeat the process for $B=10,000$ times; each time, we conduct the hypothesis test (4) at level $\alpha=0.05$, with $\widehat{\tau}_{N}$ calculated with true and estimated propensity scores, respectively. We then estimate the power by the proportion of the $B=10,000$ replications where the null hypothesis is rejected. Table 1 also reports the power calculated based on Shook-Sa and Hudgens, (2022). The proposed method consistently leads to powers close to the nominal 0.80; the method of Shook-Sa and Hudgens, (2022) leads to similar power. As noted before, the power based on the estimated propensity scores is higher than that based on the true scores, rendering the proposed method conservative. Note that the discrepancy in variance between the true and estimated propensity scores could be magnified in the simulation compared to most real studies, because both $Z$ and $Y$ are drawn from the underlying true model. We also include the sample size calculated by two-sample $z$-test for comparison. When the overlap is perfect, the $z$-test correctly calculates the sample size. However, when the overlap is poor, meaning that the intended observational study is far from a target randomized trial, the $z$-test, which treats observational studies as randomized trials and thus ignore the degree of overlap, tends to vastly underestimate the sample size.
The parameter $\rho^{2}$ is usually unknown; we suggest treating $\rho^{2}$ as a sensitivity parameter between 00 and $R^{2}$. For $\phi\in\{1.00,0.98,0.93,0.88,0.84,0.81\}$, sensitivity analyses will yield sample sizes of $(629,630)$, $(653,669)$, $(742,821)$, $(959,1239)$, $(1225,1818)$ and $(1515,2516)$, respectively. The results show that when the overlap is good, i.e., the observational study resembles a randomized trial, the minimal sample size is robust to the choice of $\rho^{2}$. However, as the overlap deteriorates, the sample size becomes more sensitive to the choice of $\rho^{2}$. This finding restates the importance of covariate overlap in observational studies, and presents the challenge of sample size calculation with worse overlap. The wide intervals do not reflect a drawback of the proposed method; instead, it displays the sensitivity to the “hidden” parameter $\rho^{2}$, which has been mostly ignored in the existing literature. The sensitivity also shows the importance of the bound $\rho^{2}\leq R^{2}$, because $R^{2}$ is often small (e.g. 0.2) in clinical studies, which guarantees that the sample size falls into a reasonable range with modest overlap.
We also conduct simulation studies where (i) the outcomes are binary and (ii) the estimand is the average treatment effect for the treated or the overlap population, the details of which are relegated to Supplementary Material D.3 and D.4, respectively. The proposed method is shown to correctly estimate the sample size under both settings.
4.2 Real data based simulations